On January 1, 2018, the Allegheny Corporation purchased machinery for $162,000. The estimated service life of the machinery is 10 years and the estimated residual value is $8,000. The machine is expected to produce 350,000 units during its life.

Required:
Calculate depreciation for 2018 and 2019 using each of the following methods.

1. Straight line.
2. Sum-of-the-years'-digits.
3. Double-declining balance.
4. One hundred fifty percent declining balance.
5. Units of production (units produced in 2018, 47,000; units produced in 2019, 42,000).
Answer #1

1. Straight line method

= (Cost - salvage value) /Years

= ($162,000 - $8,000)/10

= $15,400 per year

2018 depreciation = $15,400

2019 depreciation = $15,400

2. Sum-of-the-years-digits.

Depreciation ($162,000-$8,000) = $154,000

Year Depreciation Remaining life of machine Depreciation fraction Depreciation expense
2018 $154,000 10 10/55 $28,000
2019 $154,000 9 9/55 $25,200

Depreciation expense(2018) = ($154,000*10/55)

= $28,000

2019 = ($154,000*9/55)

= $25,200

3. Double - declining balance

Under DDB, rate is double as compare to the rate under straight line method.

Rate (SLM) = 100/10 = 10%.

Rate (DDB) = 20%

2018 depreciation( cost *rate) = $162,000*20% = $32,400

2019 = ($162,000-$32,400)*20% = $25,920

4. One hundred fifty percent declining balance

Rate = 150% of straight line method

Rate = 1.5*10%

= 15%

2018 = $162,000*15% = $24,300

2019 = ($162,000-$24,300)*15% = $20,655

5. Units of production

= [Depreciation *(units produced in a year/Total units)]

2018 = ($154,000*47,000/350,000)

= $20,680

2019 = ($154,000*42,000/350,000)

= $18,480